OpenStudy (anonymous):

Find the slope of the line through the following pair of points. (4,-2), (7,-1) Indicate the rise and run from the leftmost point to the rightmost point. rise run

OpenStudy (anonymous):

(-2-(-1))/(4-7) just reduce that and see what you get

OpenStudy (anonymous):

what do u mean reduce it , i know what reduce means but im still confused

OpenStudy (haseeb96):

m=1\3 so it indicate rise

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

by reduce I mean solve the top and bottom -2-(-1)=-1 4-7=-3 -1/-3=1/3